OpenStudy (anonymous):

The given equation is either linear or equivalent to a linear equation. Solve the equation. (If there is no solution, enter NO SOLUTION. If all real numbers are solutions, enter REALS.) (t − 7)^2 = (t + 7)^2 + 112

OpenStudy (dumbcow):

t = -4 I'll let Harkirat show solution

OpenStudy (anonymous):

using (a-b)^2 = a^2 - 2ab + b^2 and (a+b)^2 = a^2 + 2ab + b^2 t^2 - 14t + 49 = t^2 + 14t + 49 + 112 t^2 - t^2 - 14t - 14t = 49+112-49 -28t = 112 t = 112/-28 t = -4

OpenStudy (anonymous):

thanks guys!

OpenStudy (anonymous):

hey friend, click the GOOD ANSWER button for us .....☺